In a simple linear regression based on 22 observations, it is found that SSE = 2,852 and SST = 10,171.

a. Calculate s2ese2 and se. (Round your answers to 2 decimal places.)

Se2
se


b. Calculate the coefficient of determination R2. (Round your answer to 4 decimal places.)
Answer #1

Answer:

a)

Given,

To determine standard error

standard error = sqrt(SSE/(n - 2))

substitute the given values

= sqrt(2852 / (22 - 2))

= sqrt(2852/20)

Standard error= 11.94

B)

Now ,

To determine the coefficient of determination

consider,

Coefficient of determination = 1- (SSE / SST)

substitute the given values

= 1- ( 2852 / 10171)

= 1 - 0.28041

Coefficient of determination = 0.7196